5 fl. oz of a 2% alcohol solution was mixed with 11 fl oz. of a 66% alcohol solution. Find the concentration of the new mixture.
Answer by checkley75(3666)

5*.02+11*.66=16x
.1+7.26=16x
16x=7.37
x=7.37/16
x=.46 or 46% is the new mix.